OpenStudy (anonymous):

The point P is on the unit circle. Find the y-coordinate of P if the x-coordinate is (-2/5) and P is above the x-axis. A) 1/5 B) (radical 6)/5 C) (radical 21)/5 D) 3(radical 2)/5 E) 2/5

OpenStudy (amistre64):

sqrt(1-(-2/5)^2)

OpenStudy (amistre64):

sin(cos^-1(-2/5)) ?

OpenStudy (amistre64):

sqrt(1 - 4/25) = sqrt(21/25) = sqrt(21)/5 c